We started a vacuum work in this morning. And still it's going on.
Although the last night the green team replaced a steering mirror by an 80% reflector on the PLS table, the beam axis to the MC looks fine.
The MC refl beam successfully goes into the MCrefl PD, and we can see the MC flashing as usual.
We started measuring the distance of the optics inside the vacuum chamber, found the distance from MC3 to MMT1(curved mirror) is ~13cm shorter than the design.
We moved the positions of the flat mirror after the Faraday and the MMT1, but could not track the beam very well because we did not completely lock the MC.
Now we are trying to get the lock of the MC by steering the MC mirrors.
Kevin suceeded in locking it !!